The ratio of robins to blackbirds that visit our feeder is 4 to 7. If there are a total 132 birds, how many robins visit the feeder? (show your work)
Answer by fractalier(2101)

Well to do it strictly "by the book," here is what you do...
First, let b be blackbirds and r be robins.
Set up a proportion next, that is
r/b = 4/7
Now cross-multiply the proportion to get
7r = 4b
We also know that r + b = 132.
Thus b = 132 - r
Now substitute this into the above equation and get
7r = 4(132 - r)
7r = 528 - 4r
11r = 528
r = 48
thus b = 84
and we're done...